OpenStudy (cenaida):

3^3 x 3^2 = A. 3^5 B. 3^6 C. 9^5 D. 9^6

jimthompson5910 (jim_thompson5910):

Add the exponents \[\Large 3^3 \times 3^2 = 3^{3+2} = 3^5\] So, \[\Large 3^3 \times 3^2 = 3^5\]

OpenStudy (cenaida):

thanks @jim_thompson5910 !

jimthompson5910 (jim_thompson5910):

you're welcome